OpenStudy (anonymous):

Functions, Problem: -2 goes in, -8 comes out.

OpenStudy (anonymous):

Could there be multiple answers?

OpenStudy (anonymous):

Just like before. f(-2) = -8.

OpenStudy (anonymous):

yup!! i see infinite solution for this one.

OpenStudy (anonymous):

So, what do you suppose happened to the -2? There are infinite answers, but you just need to find one answer.

OpenStudy (anonymous):

x^2 - 12? Because if -2 is x then that would give you an 4, and then -12 to give you an -8?

OpenStudy (anonymous):

All you have to do is stick with the -2 and the -8. There are at least 2 easy functions that can be applied to -2 to get it to -8. One is multiplication by a number and another is addition by a number.

OpenStudy (anonymous):

So its like factoring in a way? -2 * 4?

OpenStudy (anonymous):

Think of it as "something happening" to -2 and then you get -8. You just have to find a suitable "something happening". There are many many solutions, but just make up an easy one. Yes! You're answer of multiplying by 2 is good!

OpenStudy (anonymous):

Another one could have been f(x) = y = x - 6.

OpenStudy (anonymous):

But I don't understand how you would insert that in as x? would it just be -2 * 4? Or would it have an x somewhere in the equation?

OpenStudy (anonymous):

Oh, Wait! The number going in is the x?

OpenStudy (anonymous):

You have to generalize at this point. Since (-2) * 4 = 8, then you can choose other values for x like 9. Then 9 * 4 = 36.

OpenStudy (anonymous):

So, you are trying to describe a function where you can take any value of x. So, x * 4 = "some number". So, your function is 4 * x. So, f(x) = y = 4 * x.

OpenStudy (anonymous):

Ohh! I got it now. That makes sense. I didn't realize that the number going in was x, that wasn't clicking.

OpenStudy (anonymous):

Yes, that's the conceptual part. One takes x and then "does something" with it.

OpenStudy (anonymous):

I see it now, thank you so much for your help.

OpenStudy (anonymous):

You're welcome